What is the weight in pounds of a tank that is a vertical cylinder with diameter of 6ft and height of 15ft


1
Expert's answer
2020-11-16T07:49:29-0500

The tank itself weighs 20 lbs per ft of height. We know that 1 cu/ft of water weighs 62.3 lb at 70 degrees.


"W=W_t+W_W\\W=20(15)+62.3(0.25\pi (6)^2(15))=26722.4\ lb"